(S.0.3) Larsen Company has a defined benefit pension plan which covers 220 employees. Based upon negotiations with the employees, Larsen Company amends its pension plan as of 1/1/17 and grants $111,600 of prior service cost to its employees. The following schedule reflects employee groups based on expected years of retirement. Group A B C D E Number of Employees 30 40 35 70 45 220 Expected Retirement on 12/31 2017 2018 2019 2020 2021 . Instructions: Compute the annual amortization of the prior service cost using the 1) years­ of ­service method, and 2)straight­line amortization over the average remaining service life of employees.
Answer #1

Question 1.

Years of service Method:-

Group No. of Employees Expected Retirement date Remaining Years Total Remaining Years for all employees Calculation Amount to be amortized
A 30 31-12-2017 1 30 =(220/720)*111600 $34100
B 40 31-12-2018 2 80 =190/720*111600 $29450
C 35 31-12-2019 3 105 =150/720*111600 $23250
D 70 31-12-2020 4 280 =115/720*111600 $17825
E 45 31-12-2021 5 225 =45/720*111600 $6975
Total 220 Total Service Years 15 720 $111600

As per years of service method, If the employees are going to retire according to a fixed pattern, Each year receives its own amortization fraction. The numerator will the number of employees who worked this year; the denominator (which represents the total service period) is determined under the following formula:

Denominator = [(n x (n + 1))/2] x d

d represents the decrease in employees each year, n represents the number of physical years remaining from the point the prior service costs are created until the last of the employees affected by these service costs retires.

Question 2

Straight line amortization over the average remaining service life of employees

Group No. of Employees Expected Retirement date Remaining Years Total Remaining Years for all employees
A 30 31-12-2017 1 30
B 40 31-12-2018 2 80
C 35 31-12-2019 3 105
D 70 31-12-2020 4 280
E 45 31-12-2021 5 225
Total 220 Total Service Years 15 720

If we use straight-line amortization, we divide the $111,600 by the average number of years remaining. The average is determined by dividing the total remaining years of all employees (720) by the number of employees (220), yielding 3.272727273,

= $111,600/3.272727273 = $34,100 per year

Thanks and Regards